Ganesh was traveled 40mphr in 2hrs and again he traveled 80mphr in 3 hrs.. Find
the speed of his entire journey

Answer by stanbon(75887)

Ganesh was traveled 40mphr in 2hrs and again he traveled 80mphr in 3 hrs.. Find
the speed of his entire journey
-----------------
Average speed = (total distance)/(total time)
--------
= (2*40 + 3*80)/(2+3) = 320/5 = 64 mph
-----------
Cheers,
Stan H.
-----------------